# Video Understanding with Large Language Models: A Survey

### v3

@@ II-B Background for LLMs
Language models are used to model language by calculating a joint probability distribution $p(x_{1:L})$ over sequences of tokens $x_{1:L}$, which can be broken down into a product of the conditional probabilities for each token with the chain rule:
@@
Large language models usually refer to those language models with thousands of millions (even billions) of parameters. The architecture of the LLM incorporates a text tokenizer and multiple self-attention layers as a transformer decoder. LLMs receive context prompts and pre-generated output to predict the next token probability in the autoregressive paradigm:
@@
where $\mathcal{M}$ represents the large language model.
Decoding strategies dictate how to harness the next token probability and select the next token $y_{t}$ from the set $S$ of all possible tokens in the vocabulary, which includes special tokens such as $<$SOS$>$, $<$EOS$>$, and $<$PAD$>$. Greedy decoding, the simplest strategy, selects the token with the highest probability, formalized as
@@
Besides deterministic strategies, sampling strategies that randomly select the next tokens using model probability are also popular in real applications, providing diverse outputs and enabling self-consistence methods.
Large language models (LLM) possess extensive generalization abilities that can be applied to various downstream tasks, including multi-modal tasks. Multimodel Large Language Models (MLLMs) [143, 144, 145, 146, 147] typically incorporate multi-modal encoders, cross-modality aligners, and an LLMs core structure. By combination of multi-modal encoders and the LLM core, MLLMs excel at integrating visual and linguistic contexts to produce detailed content.
@@
  • Emergent Abilities: When the parameter size and volume of training data from a large language model exceed a certain magnitude, some novel capabilities emerge, such as in-context learning, instruction following, and step-by-step reasoning [75].
@@ III Vid-LLMs
Figure 4: Taxonomy of Video Understanding with Large Language Models (Vid-LLMs), consists of Video Analyzer $\times$ LLM, Video Embedder $\times$ LLM and (Analyzer + Embedder) $\times$ LLM, and the sub-categories are LLM as Summarizer/Manger/Text Decoder/Regressor/Hidden Layer. Font color indicates the granularity of video understanding supported by the Vid-LLMs: black for abstract understanding, light red for temporal understanding, and blue for spatiotemporal understanding.
Figure 5: The figure illustrates three primary frameworks for Vid-LLMs: (1) Video Analyzer $\times$ LLM, where video analyzers convert video inputs to textual analysis for the LLM; (2) Video Embedder $\times$ LLM, where video embedders generate vector representations (embeddings) for the LLM to process; (3) (Analyzer + Embedder) $\times$ LLM, a hybrid approach combining both analyzers and embedders to provide the LLM with textual analysis and embeddings.
In this section, we introduce the taxonomy of Vid-LLMs shown as Figure 4, and the training strategies shown as Figure 6.
@@ III-A Taxonomy
Based on the method of processing input video, we classify Vid-LLMs into three primary categories: Video Analyzer $\times$ LLM, Video Embedder $\times$ LLM, and (Analyzer + Embedder) $\times$ LLM, as illustrated in Figure 5.
@@ III-A1 Video Analyzer $\times$ LLM
  • LLM as Manager: In this subcategory, the LLM primarily coordinates the overall system’s operation. It may actively generate commands to invoke various Video Analyzers to produce the desired results for the user, call the Video Analyzer and further process the obtained analysis before returning it to the user, or engage in multiple rounds of interaction with the Video Analyzer. Compared to the LLM as Summarizer category, the LLM as Manager category is more flexible and can be distinguished by its information flow complexity. Examples of Vid-LLMs in the Video Analyzer $\times$ LLM as Manager category include: ViperGPT [183], Video ChatCaptioner [271], ChatVideo [182], AssistGPT [181], HuggingGPT [171], ProViQ [180], LifelongMemory [179], DoraemonGPT [177], KEPP [176], VURF [175], VideoAgent (Stanford) [174], VideoAgent (PKU) [173], TV-TREES [172], SCHEMA [170], RAVA [169], GPTSee [165], TraveLER [168], LAVAD [167], VideoTree [166], LVNet [164], OmAgent [163], DrVideo [162], etc.
@@ III-B1 Training-free Vid-LLMs
Many Vid-LLMs systems are built on powerful LLMs with strong zero-shot, in-context learning, and Chain-of-Thought capabilities. These systems do not require training in the parameters of the LLM or other modules. Most Vid-LLMs in the Video Analyzer $\times$ LLM category are training-free because the information from the video and other accompanying modalities has already been parsed into text. At this point, the video understanding task has been transformed into a text understanding task. Since LLMs can unify almost all NLP tasks into generation tasks, they can also handle many video understanding tasks.
@@ III-B2 Fine-tuning Vid-LLMs
  • Connective Adapter Fine-tuning: Here, the term ”Connective” refers to adapters that bridge the Video Embedder and the LLM externally, enabling information from the video to flow into the LLM through the Connective Adapter. As illustrated in Figure 6, during training, the parameters of both the Video Embedder and the LLM are frozen, and only the parameters of the Connective Adapter are updated. Common Connective Adapters include MLP/Linear Layer and Q-former [276], their combinations, etc., whose primary function is to map video embeddings from the visual semantic space to the text semantic space of the LLM input tokens (i.e., modality alignment). Typically, fine-tuning only the Connective Adapter does not alter the LLM’s inherent behavior.
@@ VI Future Directions and Conclusion
In this survey, we have reviewed the latest advancements in video understanding with large language models (Vid-LLMs) and introduced the fundamental principles, significant discoveries, and techniques for understanding and applying Vid-LLMs effectively. We begin from the developmental history of video understanding, from traditional non-deep learning methods to neural network-based approaches, then to self-supervised pre-training for videos, and now to the current video understanding solutions based on LLMs.
@@ VI-A Limitations and Future Work
Although current methods enhance the effectiveness of various downstream tasks by pre-training and fine-tuning with a large amount of video data, and the introduction of LLMs allows for a better understanding of various information in videos, there remain many challenges unresolved when facing diverse video understanding tasks in the real world.
@@
  • Multi-modal Video Understanding: Multi-modal video understanding requires the integration of different types of data such as visual, audio, and text for a better understanding of videos. Aligning these data, especially in terms of spatial and temporal synchronization, is particularly critical. This area lacks relevant research and suffers from a scarcity of datasets. Moreover, constructing such datasets faces challenges as ensuring high quality and consistency in data annotation is often difficult. In terms of methodology, extracting and utilizing effective features across different modalities is key to achieving precise video understanding, but this process is filled with challenges.
@@
  • Hallucination in Multimodal LLMs: “Hallucination” refers to the phenomenon where the model generates responses that are significantly disconnected from the relevant source material or input. This can lead to the creation of highly erroneous or unrealistic descriptions that don’t align with the provided videos. In video understanding with LLMs, the main reasons for this situation are as follows: 1. Insufficient extraction of visual features. 2. The influence of video contextual content. 3. The domain gap between the visual feature domain and the Language domain. 4. Hallucination inherent in LLMs. Therefore, to address the impact of hallucination, solutions can be found in improving the effectiveness of video encoders, enhancing the understanding of long-form spatio-temporal contexts, and the collaboration between visual latent space and linguistic latent space.

### v4

@@ II-B Background for LLMs
Language models are trained to learn a joint probability distribution $p(x_{1:L})$ with a sequence of text tokens $x_{1:L}$. This joint distribution is usually equivalent to a product of the conditional probabilities conditioned on each token with the chain rule:
@@
Large Language Models (LLMs) refer to language models with a large number of parameters, e.g., billions. The architecture of LLMs incorporates a text tokenizer and multiple self-attention layers. LLMs are trained in a teacher-forcing manner to predict the next token’s probability, where the generation process utilizes the autoregressive paradigm:
@@
where $\mathcal{M}$ represents an LLM.
Decoding strategies dictate how to harness the next token probability and select the next token $y_{t}$ from the set $S$ of all possible tokens in the vocabulary, which includes special tokens such as $<$SOS$>$, $<$EOS$>$, and $<$PAD$>$. Greedy decoding, the simplest strategy, selects the token with the highest probability, formalized as:
@@
Besides deterministic strategies, sampling strategies that randomly select the next tokens using model probability are also popular in real applications. These strategies provide diverse outputs and enable self-consistency methods.
@@
  • Emergent Abilities [77]: When the parameter size and volume of training data for a large language model exceed a certain magnitude, some novel capabilities emerge, such as in-context learning, instruction following, and step-by-step reasoning. In-context learning allows the model to learn and make predictions based on the context provided within the input text without requiring explicit retraining. Instruction following enables the model to perform tasks based on natural language instructions. Step-by-step reasoning, e.g. chain-of-thought (CoT), allows the model to follow a logical sequence of steps to arrive at a conclusion, which is particularly useful for solving complex problems.
LLMs possess extensive generalization abilities that can be applied to various downstream tasks, including multi-modal tasks. Multimodal Large Language Models (MLLMs) [146, 76, 147, 148, 149] typically incorporate multimodal encoders, cross-modality aligners, and an LLM core structure. By combining multimodal encoders with the LLM core, MLLMs excel at integrating visual and linguistic contexts to produce detailed content.
@@ III Vid-LLMs
In this section, we introduce a novel taxonomy of Vid-LLMs, providing a comprehensive overview of their classification. Following this, we explore the diverse training strategies that empower Vid-LLMs to achieve their capabilities.
Figure 4: The figure illustrates three primary frameworks for Vid-LLMs: (1) Video Analyzer $\times$ LLM, where video analyzers convert video inputs to textual analysis for the LLM; (2) Video Embedder $\times$ LLM, where video embedders generate vector representations (embeddings) for the LLM to process; (3) (Analyzer + Embedder) $\times$ LLM, a hybrid approach combining both analyzers and embedders to provide the LLM with textual analysis and embeddings. The arrows indicate the direction of information flow, with dashed arrows representing optional paths. Blue arrows denote textual information flow, while red arrows denote embeddings.
@@
Figure 5: Taxonomy of Video Understanding with Large Language Models (Vid-LLMs), consists of Video Analyzer $\times$ LLM, Video Embedder $\times$ LLM and (Analyzer + Embedder) $\times$ LLM, and the sub-categories are LLM as Summarizer/Manager/Text Decoder/Regressor/Hidden Layer. Font color indicates the granularity of video understanding supported by the Vid-LLMs: black for abstract understanding, red for temporal understanding, and blue for spatiotemporal understanding.
@@ III-A Taxonomy
Based on the method of processing input video, we categorize Vid-LLMs into three primary types: Video Analyzer $\times$ LLM, Video Embedder $\times$ LLM, and (Analyzer + Embedder) $\times$ LLM. Each category represents a unique approach to integrating video processing with LLMs, as illustrated in Figure 4.
@@ III-A1 Video Analyzer $\times$ LLM
  • LLM as Manager: In this subcategory, the LLM primarily coordinates the overall system’s operation. It may actively generate commands to invoke various Video Analyzers to produce the desired results for the user, call the Video Analyzer and further process the obtained analysis before returning it to the user, or engage in multiple rounds of interaction with the Video Analyzer. Compared to the LLM as Summarizer category, the LLM as Manager category is more flexible and can be distinguished by its information flow complexity. Examples of Vid-LLMs in the Video Analyzer $\times$ LLM as Manager category include: ViperGPT [184], Video ChatCaptioner [272], ChatVideo [183], AssistGPT [182], HuggingGPT [172], Hawk [179], ProViQ [181], LifelongMemory [180], DoraemonGPT [178], KEPP [177], VURF [176], VideoAgent (by Stanford) [175], VideoAgent (by PKU) [174], TV-TREES [173], SCHEMA [171], RAVA [170], GPTSee [166], TraveLER [169], LAVAD [168], VideoTree [167], LVNet [165], OmAgent [164], DrVideo [163], etc.
@@ III-B1 Training-free Vid-LLMs
Many Vid-LLMs systems are built on powerful LLMs with strong zero-shot, in-context learning, and Chain-of-Thought capabilities. These systems do not require training in the parameters of the LLM or other modules. Most Vid-LLMs in the Video Analyzer $\times$ LLM category are training-free because the information from the video and other accompanying modalities has already been parsed into text. At this point, the video understanding task has been transformed into a text understanding task. Since LLMs can unify almost all NLP tasks into generation tasks, they can also handle many video understanding tasks. SlowFast-LLaVA [273] is a training-free Video LLM that uses a two-stream input design to capture both spatial semantics and temporal context without fine-tuning and demonstrates capabilities across various video understanding benchmarks.
@@ III-B2 Fine-tuning Vid-LLMs
  • Connective Adapter Fine-tuning: Here, the term “Connective” refers to adapters that bridge the Video Embedder and the LLM externally, enabling information from the video to flow into the LLM through the Connective Adapter. As illustrated in Figure 6, during training, the parameters of both the Video Embedder and the LLM are frozen, and only the parameters of the Connective Adapter are updated. Common Connective Adapters include MLP/Linear Layer and Q-former [278], their combinations, etc., whose primary function is to map video embeddings from the visual semantic space to the text semantic space of the LLM input tokens (i.e., modality alignment). Typically, fine-tuning only the Connective Adapter does not alter the LLM’s inherent behavior.
@@ VI Future Directions and Conclusion
In this survey, we have reviewed the latest advancements in video understanding methods with large language models (Vid-LLMs) and introduced the fundamental principles, significant discoveries, and techniques for understanding and applying Vid-LLMs effectively. We begin with the developmental history of video understanding, from traditional non-deep learning methods to neural network-based approaches, then self-supervised pretraining for videos, and now the current video understanding solutions based on LLMs.
@@ VI-A Limitations and Future Work
Although current methods enhance the effectiveness of various downstream tasks by pretraining and fine-tuning with a large amount of video data, and the introduction of LLMs allows for a better understanding of various information in videos, there remain many challenges unresolved when facing diverse video understanding tasks in the real world.
@@
  • Multimodal Video Understanding: Multimodal video understanding requires the integration of different types of data, such as visual, audio, and text, for a better understanding of videos. Aligning these data, especially in terms of spatial and temporal synchronization, is particularly critical. This area lacks relevant research and suffers from a scarcity of datasets. Moreover, constructing such datasets faces challenges as ensuring high quality and consistency in data annotation is often difficult. In terms of methodology, extracting and utilizing effective features across different modalities is key to achieving precise video understanding, but this process is filled with challenges.
@@
  • Hallucination in Multimodal LLMs: “Hallucination” refers to the phenomenon where the model generates responses that are significantly disconnected from the relevant source material or input. This can lead to the creation of highly erroneous or unrealistic descriptions that do not align with the provided videos. In video understanding with LLMs, the main reasons for this situation are as follows: 1) Insufficient extraction of visual features. 2) The influence of video contextual content. 3) The domain gap between the visual feature domain and the Language domain. 4) Hallucination inherent in LLMs. Therefore, to address the impact of hallucination, solutions can be found in improving the effectiveness of video encoders, enhancing the understanding of long-form spatio-temporal contexts, and the collaboration between visual latent space and linguistic latent space.